# DiffRegCD: Integrated Registration and Change Detection with Diffusion Features

### v1

@@ Abstract
Change detection (CD) is critical in computer vision and remote sensing, with applications in monitoring, disaster response, and urban analysis. Most CD models assume co-registered inputs, but real imagery often suffers from parallax, viewpoint shifts, or long temporal gaps, leading to severe misalignment. Conventional register-then-detect pipelines and recent joint frameworks (e.g., BiFA, ChangeRD) remain limited: they rely on regression-only flow, global homographies, or synthetic perturbations that fail under large displacements. We propose DiffRegCD, an integrated framework that couples dense registration and change detection. DiffRegCD reformulates correspondence as a Gaussian-smoothed classification task, delivering sub-pixel accuracy and stable training. It builds on frozen multi-scale features from a pretrained denoising diffusion model, which provide invariance to viewpoint and illumination variation. Supervision is enabled by controlled affine perturbations applied to standard CD datasets, yielding paired ground truth for both flow and change detection without pseudo-labels. Experiments on aerial (LEVIR-CD, DSIFN-CD, WHU-CD, SYSU-CD) and ground-level (VL-CMU-CD) datasets show that DiffRegCD outperforms recent baselines and remains robust under wide temporal and viewpoint variation, establishing diffusion features and classification-based correspondence as a strong foundation for integrated CD. Code will be released after the review process.
@@ 1 Introduction
Training is enabled through controlled affine perturbations applied to existing CD datasets (e.g., VL-CMU-CD [alcantarilla2016street]), yielding paired ground truth for both flow and change labels. This principled strategy avoids pseudo-labels and handcrafted augmentation. Once trained, DiffRegCD generalizes across ground-level and aerial datasets, including VL-CMU-CD, ChangeSim [park2021changesimendtoendonlinescene], LEVIR-CD, DSIFN-CD, WHU-CD, and SYSU-CD, demonstrating robustness to long temporal gaps and wide viewpoint variation.
@@ 2.4 Discussion
In summary, registration methods range from sparse keypoints to dense transformers, but all degrade under long temporal baselines. CD models achieve high performance in curated, co-registered benchmarks but collapse when misalignment is present. Joint frameworks attempt to address this gap, yet are constrained by implicit alignment strategies, regression instability, or unrealistic training signals.
Our work differs by explicitly integrating dense registration and change detection. DiffRegCD leverages diffusion-pretrained features for viewpoint- and illumination-invariant representations, reformulates correspondence as a classification task for sub-pixel alignment, and introduces supervised multi-scale warping with paired flow and change labels. This design enables robust change detection under severe misalignment, bridging the gap between registration accuracy and semantic consistency.
@@ 3.1 Diffusion-Pretrained Features
We employ a frozen denoising diffusion encoder $\phi$, trained on large-scale natural imagery, as a universal feature extractor. For an input $I$, $\phi$ provides a hierarchy of features indexed both by timestep $t\!\in\!\mathcal{T}$ (diffusion noise level) and scale $i\!\in\!0{:}L$ (resolution in the encoder pyramid):
@@
Small $t$ correspond to lightly noised inputs and thus preserve high-frequency edges and textures; large $t$ are closer to semantic denoising stages, encoding shape and category information invariant to illumination, season, or sensor modality. This duality provides both local detail and global invariance, which are critical for misaligned change detection. The spatial pyramid captures features from coarse (contextual, semantic) to fine (local, geometric) levels. Together with timestep variation, this yields a two-dimensional grid of features $\{f^{t,i}\}$ spanning both resolution and semantic abstraction. Since $C_{i}$ varies with scale and timestep, each feature map is projected to a fixed dimension $D_{i}$ with a $1{\times}1$ linear map $P_{i}$ followed by LayerNorm:
@@ 3.2 Gaussian-Processed Transformer for Flow
Gaussian pre-smoothing. We estimate flow from the coarse semantic DDPM[ho2020denoisingdiffusionprobabilisticmodels] feature at block $i^{\star}{=}11$ and timestep $t^{\star}$. Before entering the transformer, $(\bar{f}_{A}^{t^{\star},11},\bar{f}_{B}^{t^{\star},11})$ are smoothed by an isotropic Gaussian kernel $\mathcal{K}_{\sigma}$ (CosKernel):
$\tilde{f}^{t^{\star},11}(x)=\sum_{p\in\mathcal{N}(x)}\exp\!\left(-\tfrac{\|x-p\|^{2}}{2\sigma^{2}}\right)\bar{f}^{t^{\star},11}(p).$
This operation encodes a Gaussian Process prior: nearby positions are correlated, uncertainty grows with distance, and the decoder inherits a smooth inductive bias.
Transformer decoder. A cross-attention stack $\Psi$ of $S$ blocks is applied to smoothed features. Queries come from $\tilde{f}_{A}^{t^{\star},11}$, keys/values from $\tilde{f}_{B}^{t^{\star},11}$:
$Q=W_{Q}(\tilde{f}_{A}^{t^{\star},11}{+}\pi),\;\;K=W_{K}(\tilde{f}_{B}^{t^{\star},11}{+}\pi),\;\;V=W_{V}(\tilde{f}_{B}^{t^{\star},11}),$
with $\mathrm{Attn}(Q,K,V)=\mathrm{softmax}(QK^{\top}/\sqrt{d})V$. The output $h^{t^{\star},11}$ encodes coarse correspondences and feeds the registration head.
@@ 3.4 Curriculum and Joint Optimization
Figure 2: Cross-dataset qualitative results under induced misregistration. Each mini-panel follows the same layout: left—inputs $I_{t}$ and $I_{t+1}$; right—top shows ground-truth flow (color-wheel) and change mask; bottom shows our predicted flow and change map. We display diverse aerial and street-level scenes from LEVIR-CD, WHU-CD, DSIFN, SYSU-CD, and VL-CMU-CD. Patch size $256{\times}256$; level Hard: $\Delta x,y\!\in\![-25,25]$ px, $\theta\!\in\![-30^{\circ},30^{\circ}]$, $s\!\in\![0.80,1.25]$. Beyond alignment, our method yields crisp boundaries, fewer background false alarms, and better recovery of small structures across datasets. White denotes change; threshold 0.5; no post-processing.
@@ 4.3 Robustness to Induced Misalignment
Figures 32 illustrate how our framework improves both registration and change localization.
@@ 4.4 Qualitative Results
LEVER-CD (Ablation Variants) mF1 mIoU OA F11
@@
LEVER-CD (External Baselines) mF1 mIoU OA F11
@@ 5 Conclusion
Limitations and Future Work. The current framework remains memory-intensive due to multi-scale aggregation and does not explicitly capture uncertainty in flow prediction. Future directions include adaptive resolution strategies, uncertainty-aware flow decoding, and incorporating additional modalities (e.g., LiDAR, hyperspectral) to further enhance robustness for long-term, real-world monitoring.

### v3

@@ Abstract
Change detection (CD) is critical in computer vision and remote sensing, with applications in monitoring, disaster response, and urban analysis. Most CD models assume co-registered inputs, but real imagery often suffers from parallax, viewpoint shifts, or long temporal gaps, leading to severe misalignment. Conventional register-then-detect pipelines and recent joint frameworks (e.g., BiFA, ChangeRD) remain limited: they rely on regression-only flow, global homographies, or synthetic perturbations that fail under large displacements. We propose DiffRegCD, an integrated framework that couples dense registration and change detection. DiffRegCD reformulates correspondence as a Gaussian-smoothed classification task, delivering sub-pixel accuracy and stable training. It builds on frozen multi-scale features from a pretrained denoising diffusion model, which provide invariance to viewpoint and illumination variation. Supervision is enabled by controlled affine perturbations applied to standard CD datasets, yielding paired ground truth for both flow and change detection without pseudo-labels. Experiments on aerial (LEVIR-CD, DSIFN-CD, WHU-CD, SYSU-CD) and ground-level (VL-CMU-CD) datasets show that DiffRegCD outperforms recent baselines and remains robust under wide temporal and viewpoint variation, establishing diffusion features and classification-based correspondence as a strong foundation for integrated CD. The code is available at GitHub.
@@ 1 Introduction
Training is enabled through controlled affine perturbations applied to existing CD datasets (e.g., VL-CMU-CD [alcantarilla2016street]), yielding paired ground truth for both flow and change labels. This principled strategy avoids pseudo-labels and handcrafted augmentation. Once trained, DiffRegCD generalizes across ground-level and aerial datasets, including VL-CMU-CD, LEVIR-CD, DSIFN-CD, WHU-CD, and SYSU-CD, demonstrating robustness to long temporal gaps and wide viewpoint variation.
@@ 2.4 Discussion
Existing registration methods, from sparse keypoints to dense transformers, degrade under long temporal baselines, while CD models perform well only under accurate co-registration and fail under misalignment. Joint frameworks partially address this but remain limited by implicit alignment or unstable supervision. Our approach explicitly unifies dense registration and change detection. DiffRegCD leverages diffusion-pretrained, invariant features, formulates correspondence as classification for sub-pixel alignment, and applies supervised multi-scale warping with paired flow and change labels, enabling robust CD under severe misalignment.
@@ 3.1 Diffusion-Pretrained Features
We employ a frozen denoising diffusion encoder $\phi$, trained on large-scale natural imagery, as a universal feature extractor. For an input $I$, $\phi$ provides a hierarchy of features indexed by the diffusion timestep $t\in\mathcal{T}$ and the pyramid scale $i\in\{0,\dots,4\}$, corresponding to $L=4$ encoder levels spanning resolutions from full scale down to $1/16$:
@@
Small $t$ correspond to lightly noised inputs and thus preserve high-frequency edges and textures; large $t$ are closer to semantic denoising stages, encoding shape and category information invariant to illumination, season, or sensor modality. Here, “large” timesteps refer to late-stage diffusive steps with high noise levels, and we select a large but not necessarily the final timestep in practice. This duality provides both local detail and global invariance, which are critical for misaligned change detection. The spatial pyramid captures features from coarse (contextual, semantic) to fine (local, geometric) levels. Together with timestep variation, this yields a two-dimensional grid of features $\{f^{t,i}\}$ spanning both resolution and semantic abstraction. Since $C_{i}$ varies with scale and timestep, each feature map is projected to a fixed dimension $D_{i}$ with a $1{\times}1$ linear map $P_{i}$ followed by LayerNorm:
@@ 3.2 Gaussian-Processed Transformer for Flow
Fourier GP matching. We estimate flow from the coarse semantic DDPM [ho2020denoisingdiffusionprobabilisticmodels] feature at decoder block $i^{\star}{=}11$ and timestep $t^{\star}$, where $i^{\star}$ denotes the depth within the diffusion U-Net and is distinct from the multi-scale pyramid index. Before the transformer, the coarse features $(\bar{f}_{A}^{t^{\star},11},\bar{f}_{B}^{t^{\star},11})$ are processed by a GP-style matching module with a cosine kernel and Fourier feature mapping,
$k(f_{i},f_{j})\approx\phi(f_{i})^{\top}\phi(f_{j}),\qquad\phi(f)=[\cos(Wf),\sin(Wf)].$
This module produces a displacement-wise correlation volume that serves as input to the flow transformer decoder and induces a smooth GP-like prior over the predicted flow. Transformer decoder. At the selected coarse scale, we concatenate the GP posterior and image features and reshape them into a sequence of spatial tokens with positional encoding:
$T=\mathrm{reshape}([\mathrm{GP},f])+\pi,\quad T\in\mathbb{R}^{B\times HW\times C}.$
The tokens are processed by a stack of $S$ transformer blocks with self-attention,
$h^{t^{\star},11}=\Psi(T),$
followed by a linear projection that maps each token to classifier logits and an auxiliary channel. The output is reshaped back to the spatial grid and passed to the registration head for motion decoding.
@@ 3.4 Curriculum and Joint Optimization
Figure 2: Cross-dataset qualitative results under induced misregistration. Each mini-panel follows the same layout: left—inputs $I_{t}$ and $I_{t+1}$; right—top shows ground-truth flow (color-wheel) and change mask; bottom shows our predicted flow and change map. We display diverse aerial and street-level scenes from LEVIR-CD, WHU-CD, DSIFN-CD, SYSU-CD, and VL-CMU-CD. Patch size $256{\times}256$; level Hard: $\Delta x,y\!\in\![-25,25]$ px, $\theta\!\in\![-30^{\circ},30^{\circ}]$, $s\!\in\![0.80,1.25]$. Beyond alignment, our method yields crisp boundaries, fewer background false alarms, and better recovery of small structures across datasets. White denotes change; threshold 0.5; no post-processing.
@@ 4.3 Robustness to Induced Misalignment
Figures 2 – 3 illustrate how our framework improves both registration and change localization.
@@ 4.4 Qualitative Results
LEVIR-CD (Ablation Variants) mF1 mIoU OA F11
@@
LEVIR-CD (External Baselines) mF1 mIoU OA F11
@@ 5 Conclusion
Limitations and Future Work. The current framework remains memory-intensive due to multi-scale aggregation and does not explicitly model flow uncertainty. Future work includes adaptive resolutions, uncertainty-aware decoding, and incorporating additional modalities (e.g., LiDAR, hyperspectral) to enhance long-term real-world robustness.
Acknowledgements
This work was supported by the NSF CAREER award under Grant 2045489.